### v1

# Strong electron-electron interactions in a dilute weakly-localized metal near a metal-to-insulator transition

Abstract
Because it is easily switched from insulator to metal either via chemical doping or electrical gating, silicon is at the core of modern information technology and remains a candidate platform for quantum computing. The metal-to-insulator transition in this material has therefore been one of the most studied phenomena in condensed matter physics, and has been revisited with considerable profit each time a new fabrication technology has been introduced. Here we take advantage of recent advances in creating ultra-thin layers of Bohr-atom-like dopants to realize the two-dimensional disordered Hubbard model at half-filling and its metal-to-insulator transition (MIT) as a function of mean distance between atoms. We use gas-phase dosing of dopant precursor molecules on silicon to create arsenic and phosphorus $\delta$-layers as thin as 0.4 nm and as dilute as 1013 cm-2. On approaching the insulating state, the conventional weak localization effects, prevalent at high dopant densities and due to orbital motion of the electrons in the plane, become dominated by electron-electron interaction contributions which obey a paramagnetic Zeeman scaling law. The latter make a negative contribution to the conductance, and thus cannot be interpreted in terms of an emergent Kondo regime near the MIT.
The physics of two-dimensional (2D) electron gases has been the subject of extensive experimental and theoretical research efforts since the 1960s, and has led to some of the most remarkable discoveries in solid-state physics, including the quantum Hall effect cage2012quantum and topological quantum states of matter QSHI_HgTe . Despite considerable advances, 2D materials remain the subject of intense ongoing research because of the many possible consequences of combined disorder, inter-site hopping of electrons, and Coulomb interaction effects. A problem that has proven to be elusively difficult to solve is the nature of the metal-to-insulator transition (MIT) in two-dimensions. It was widely thought to be non-existent in the 80s and 90s, because of predictions that localization would turn any 2D material into an insulator PhysRevLett.42.673 , but was later observed in most 2D systems Graphene_MIT_2011 ; VdW_MIT_2019 ; SiGe_2020 ; GaAs_1999 ; AlAs_2013 ; ZnO_2006 ; Zhang:2016aa . While the 2D MIT has been studied in many materials, the mechanisms driving it and the phases associated with it remain open questions SHASHKIN2021168542 ; kirk20192d ; arovas2022hubbard . In silicon, 2D electron layers can be formed in a number of ways, for example with MOSFETs PhysRevB.50.8039 , quantum wells doi:10.1063/1.1639507 , and $\delta$-doped layers schubert1996delta . In these systems the electron mobility and the disorder strength are substantially different, as visible in Table 1; nonetheless they all exhibit the MIT at similar interaction strength parameter $r_{s}\approx 20$ PhysRevB.99.081106 , defined as the ratio of the Coulomb energy and the Fermi energy ($r_{s}=g_{\text{v}}/(\pi n_{\text{2D}})^{1/2}a_{\text{B}}$), where $n_{\text{2D}}$ is the sheet electron density, $a_{\text{B}}$ the effective Bohr radius, and $g_{\text{v}}$ the valley degeneracy of the host material). The critical value of $r_{s}$ depends on the disorder strength and is lower at high disorder Walsh_2019 . There is no agreement on the mechanisms driving the MIT dolgopolov2019 , but a number of distinct insulating phases have been proposed for clean SHASHKIN2021168542 and disordered systems Yildiz_2009 ; DasSarma_crossover ; DasSarma_perc , differing mostly in their magnetic behavior potter2012quantum .
2D system $\ell$ [nm] $\ell_{\phi}$ [nm] $n_{\text{2D}}[$1014cm-2] $\delta$ [nm] $r_{s}$
Current work 3 70 0.12-1.9 0.4-1.8 1.4-5.5
Si:As &\&& Si:P
Ge:P PhysRevLett.112.236602 20 400 0.3-1.4 1.49 0.6-1.3
Si MOSFET MC_silicon_inversion_layers 48 252 0.4-0.8$\times$10-4 10 10-20
QW 2DHG Shashkin_2005 26 61 0.4-3.8$\times$10-4 0.5 10-20
Graphene PhysRevLett.107.166602 50 900 0.005 0.5-1 $\sim$ 6-30
Table 1: Characteristics of various 2D electron gases. The systems in bold are half-filled. $\ell$ is the mean free path, $\ell_{\phi}$ the coherence length, $n_{\text{2D}}$ the 2D sheet electron density, $\delta$ the electron layer thickness and $r_{s}$ the interaction strength.
[…]
Here we report on the low temperature magneto-conductance of arsenic- and phosphorus-doped silicon $\delta$-layers at various doping densities near the MIT. Owing to the unprecedented combination of high densities and thinness for such $\delta$-layers in silicon (see Table. 1), we are able to access and highlight a new regime dominated by strong Coulomb interactions, where the interaction parameter is in the range $r_{s}=1.4-5.5$ for an unusually high density of carriers. In this system the number of free electrons corresponds to the number of donors, as confirmed by the consistency between Hall effect and X-ray fluorescence data SiAs_Xray_fluorescence . It should thus be well described by a highly disordered half-filled 2D Hubbard model. Recent theoretical studies of the ordered half-filled 2D Hubbard model Walsh_2019 predict that at low temperature and low interaction strength there is a first order transition from the metallic to the insulating phase, as was already suggested by Mott in the site-ordered Hubbard model with no disorder mott2004metal . As usual, the discontinuous transition comes with a coexistence region.
Following general arguments bergman1977critical , the addition of disorder will render the transition of percolative nature, being continuous in full equilibrium, while hysteretic behavior and a coexistence of various metastable states, that differ in their local conduction properties, is expected to be inherited from the disorder-free limit.
Figure 2: Typical magneto-conductance data. (a) At temperature $T=115{\rm}$ mK, the magnetic field is set parallel to the plane of the $\delta$-layer (blue points - with no dependence on the relative orientation to the current) and perpendicular to it (green points). The black points are the difference between in- and out-of-plane magneto-conductance per square, the solid black line being a fit with weak-localization corrections, Eq. (1) minus Eq. (2). The orange data (right y-axis) is a Hall measurement ($R_{xy}$) taken at 1.75 K, fitted linearly by the solid line. (b) Difference between out-of-plane and in-plane magneto-conductance per square for various temperatures. The solid lines are fits with Eq. (1) minus Eq. (2). (c) Conductance per square versus temperature in black, displaying a logarithmic dependence as expected for weak localization down to $T_{\text{c}}$approx 300 mK. The coherence length $\ell_{\phi}$ obtained from the fits in (b) is shown in orange. It follows a power law $T^{-0.31}$ down to $T_{\text{c}}$. All data in (a), (b), and (c) are obtained from a sample of electron density $n=(1.18\pm 0.01)\times 10^{13}$ cm-2.
Two-dimensional dopant layers in silicon
[…]
The temperature dependence of the samples’ longitudinal conductivity ($\sigma_{xx}\equiv 1/R_{xx}\times L/W$ with $L$ and $W$ the sample length and width, respectively) in the absence of magnetic field is shown in Fig. 2(c). It exhibits two regimes. The conductivity decreases logarithmically with decreasing temperature $T$ down to about $T_{\text{c}}=300$ mK, with a dependence $\sigma_{xx}\propto\ln{T}$, and saturates at lower temperatures. Such a temperature dependence has been observed in many different materials 2_to_3_crossover ; Polley:2011aa ; CHEUNG19941225 ; FOURNIER20001941 ; IHN20081851 ; PhysRevB.78.125409 , but the origin and intrinsic nature of the saturation remains debated RevModPhys.91.011002 . At temperatures above the saturation of the conductivity, the electronic system is a conventional weakly-localized two-dimensional conductor with strong electron-electron interactions.
Conventional weakly-localized regime
In the two-dimensional dopant layers studied here, the conduction at temperatures above the crossover temperature $T_{\text{c}}$ is characterized by a logarithmic temperature dependence of the conductivity which remains larger than the quantum of conductance $G_{0}=e^{2}/h$, indicative of diffusive transport, for which weak-localization (WL) effects can be expected PhysRevB.22.5142 . Weak localization is a quantum correction to the Drude conductivity that stems from the interference of diffusive electrons with themselves through self-intercepting loops. For weak localization to arise the electron coherence length $\ell_{\phi}$ must exceed the mean free path $\ell$, which sets the scale of the smallest loops. Weak localization is strongest if time reversal symmetry is not broken; when a magnetic field is applied, the time-reversal symmetry is broken as the field couples to the electron’s orbital motion. For a two-dimensional conduction layer with the field perpendicular to the layer $B_{\perp}$, the ensuing change in conductivity is given by the following expression derived by Hikami, Larkin and Nagaoka HLN :
$\Delta\sigma(B_{\perp})=\frac{G_{0}}{\pi}\left[\psi\left(\frac{1}{2}+\frac{B_{\phi}}{B_{\perp}}\right)-\psi\left(\frac{1}{2}+\frac{B_{\ell}}{B_{\perp}}\right)+\ln\left(\frac{2\ell_{\phi}^{2}}{\ell^{2}}\right)\right],$ (1)
where $\psi(x)$ is the digamma function. The phase breaking field is given by $B_{\phi}=\frac{\hbar}{4e\ell_{\phi}^{2}}$ and the elastic characteristic field by $B_{\ell}=\frac{\hbar}{2e\ell^{2}}$. If the layer has a finite thickness $\delta<\ell_{\phi}$, a magnetic field applied in the conduction plane $B_{\parallel}$ still has a small orbital effect
$\Delta\sigma(B_{\parallel})=\frac{G_{0}}{\pi}\ln(1+\gamma B_{\parallel}^{2}),$ (2)
that depends on the thickness of the layer through the single parameter $\gamma=\delta^{2}\sqrt{\frac{4\pi}{n}}\left(\frac{e}{\hbar}\frac{\ell_{\phi}}{\sqrt{\ell}}\right)^{2}$ sullivan ; in_plane_MC .
Figure 3: Influence of dopant density on the magneto-conductance. (a) Magneto-conductance per square at 600 mK for samples with different free carrier densities $n_{\text{2D}}$. As $k_{\text{F}}\ell$ decreases with decreasing density the effect of the perpendicular magnetic field (full dots) becomes smaller. The conductivity for $B_{\parallel}$ is denoted by crosses. (b) Magnitude of the weak-localization part of the magneto-conductance ($\Delta\sigma_{xx}(B_{\perp})-\Delta\sigma_{xx}(B_{\parallel})$ shown on the left $y$-axis) at 600 mK in a 2 T magnetic field as a function of $k_{\text{F}}\ell$. The right $y$-axis indicates the free carrier density. The out-of-plane magneto-conductance is seen to be dominated by weak localization corrections which depend on $k_{\text{F}}\ell$. In (b) the dots correspond to arsenic $\delta$-layers while the squares are from phosphorus $\delta$-layers, showing that the magneto-conductance does not depend on the dopant species.
Electron-electron interactions in dilute $\delta$-layers
Figure 2(a) shows in-plane and out-of-plane magneto-conductance per square obtained at 115 mK. At this temperature and sample density ($n=(1.18\pm 0.01)\times 10^{13}$ cm-2) the magneto-conductance is negative. This cannot be explained by weak localization, nor can the data be fit with weak anti-localization PhysRevB.53.3912 , nor with Minkov et al.’s model which includes Rashba and Dresselhaus spin-orbit interaction PhysRevB.70.155323 ; ILP_WAL ; In_plane_InSb , as could be expected for samples with large spin-orbit coupling. Instead, the magneto-conductance corrections are governed by two effects that add up in conduction as two channels in parallel: $\Delta\sigma_{xx}^{\rm tot}=\Delta\sigma_{xx}^{\rm WL}+\sigma_{xx}^{\rm Zeeman}$ PhysRevB.26.4009 ; PhysRevLett.47.1758 . The first term $\Delta\sigma_{xx}^{\text{WL}}$ is the conventional positive weak localization as described in Eq. (1) and Eq. (2). The second term $\Delta\sigma_{xx}^{\text{Zeeman}}$ has the opposite sign. It is the Zeeman effect in the Altshuler-Aronov (AA) AAcontribution corrections to the conductivity as described in 1982 by Lee and Ramakrishnan PhysRevB.26.4009 and later elaborated on by Zala, Narozhny and Aleiner PhysRevB.64.214204 ; zala2001interaction :
$\Delta\sigma(h)=-\frac{G_{0}}{2\pi}F\ln\left(\frac{h}{1.3}\right),\text{ for }h\gg 1.$ (3)
Here $h=\frac{g\mu_{\text{B}}B}{k_{\text{B}}T}$ is the reduced magnetic field with the $g$-factor $g=2$, $\mu_{\text{B}}$ the Bohr magneton and $k_{\text{B}}$ the Boltzmann constant, and $F$ is a dimensionless number (resulting from the angular average of the statically screened Coulomb scattering). The effect scales as $\ln(B/T)$, which becomes strong at low temperatures.
The Zeeman effect results from the coupling of the magnetic field to the spin of the electrons, and thus (to the extent that spin orbit coupling does not interfere) is isotropic in the direction of the magnetic field, unlike weak localization that depends mostly on the in-plane orbital motion and thus senses the perpenicular component of the magnetic field 2_to_3_crossover . Thus, we can isolate the anisotropic WL contribution by taking the difference between the out-of-plane and in-plane magneto-conductance, as is done in Fig. 2(a) (black dots), and Fig. 2(b), where the full lines are fits with the weak-localization expression Eq. (1) from which Eq. (2) has been subtracted. In Fig. 2(c) the orange points are the coherence lengths $\ell_{\phi}$ obtained from fitting the data at different temperatures, exhibiting an effective power law ($\ell_{\phi}\sim T^{-0.31}$) that reflects the decreasing rate of inelastic scattering PhysRevB.70.245423 ; 2_to_3_crossover . Note that when fitting these curves, only the coherence length $\ell_{\phi}$ and the thickness parameter $\gamma$ are free parameters. The mean free path $\ell$, the free carrier density and mobility are all obtained from Hall measurements at high temperatures (2 K), as shown in Fig. 2(a) in orange. In particular there is no room for a free pre-factor to Eq. (1), even though it is often used with little or no justification for similar data PhysRevLett.112.236602 ; siP_sheet (though for $\sigma_{xx}\lesssim 10G_{0}/\pi$ such a prefactor may effectively arise due to two-loop localization corrections Aleiner_1999 ; Minkov_alpha ). For thicker and denser samples, the out-of-plane magneto-conductance data could (erroneously) be fitted to weak localization only, using a free pre-factor, because the AA correction is small. However, when subtracting out the AA term, no unjustified prefactor is needed. From the fit parameters we can calculate the thickness of the dopant layer sullivan , between 0.4 and 1.8 nm for the arsenic $\delta$-layers studied in this work, comparable to the silicon unit cell $\sim$0.5 nm. These are thinner than previously obtained layers in phosphorus doped Si doi:10.1063/1.4998712 ; PhysRevB.101.245419 , as is expected due to arsenic’s slower diffusion Solmi_diffusivity during annealing steps of sample fabrication, and corroborated by photoemission SiAs_ARPES and X-ray reflectometry SiAs_XRR experiments. It is interesting to note that in the thinnest layer only three Si(100) crystal planes are conductive, compared to thirteen in the thickest layer.
Figure 4: $B/T$ scaling of the AA correction. (a) Change in conductivity due to the Zeeman splitting in units of the conductance quantum $e^{2}/\hbar$, as a function of the logarithm of the reduced field $h=g\mu_{\text{B}}B/k_{\text{B}}T$. The data are shown for multiple temperatures and stem from a single sample with free carrier density $n=(2.14\pm 0.02)\times 10^{13}$ cm-2. The data collapse down to 250 mK. (b) Same as in (a), except that an effective temperature $T_{\rm eff}$ is used to obtain a full collapse of the data. $T_{\text{eff}}$ is plotted for all samples in (c). (c) Effective temperature $T_{\rm eff}$ determined for all samples so as to collapse the AA correction as in (a) and (b).
The weak localization effect calculated in Eq. (1) stems from an expansion in $(k_{\text{F}}\ell)^{-1}$, where $k_{\text{F}}$ is the Fermi momentum of the electrons, which is proportional to the square root of the dopant density. For samples with different densities, the weak localization correction depends primarily on $k_{\text{F}}\ell$, which is consistent with the monotonic dependence of the perpendicular magneto-conductance as a function of $k_{\text{F}}\ell$, as shown in Fig. 3(a) and Fig. 3(b) for $B=2~T$ (see the supplemental section I for information on the variability of $\ell$ and $k_{\text{F}}$ across samples). The electron-electron interaction effects depend on the ratio of Coulomb interaction and kinetic energy, which is proportional to the ionization energy divided by the hopping integral (the donors’ wavefunction overlap) and thus is stronger for lower carrier densities. For in-plane magnetic fields, when the $\delta$-layer is thin and/or dilute enough, the electron-electron interaction contributions dominate over the weak localization term. This is the case for our arsenic $\delta$-layers, as is visible in Fig. 3(a). This regime of thin and dilute half-filled layers has not been achieved in $\delta$-layers before.
To isolate the Zeeman terms in the Altshuler-Aronov corrections we remove the weak localization contribution to the magneto-conductance as determined by the fits shown in Fig. 2(a) and Fig. 2(b). The remaining electron-electron interaction contribution is given by Eq. (3), which is a logarithmic function of the reduced field $h=\frac{g\mu_{\text{B}}B}{k_{\text{B}}T}$, and depends on $r_{s}$ through the average Coulomb scattering parameter $F$. Therefore, when the Zeeman contribution to the magneto-conductance is plotted against the reduced field $h$ at different temperatures, the data collapses to a single curve. This is shown in Fig. 4(a) for a sample with density $n=(2.14\pm 0.02)\times 10^{13}$ cm-2, where a clear collapse of the data is seen down to 250 mK. This curve scales as $\ln(h)$ at large fields, as expected from Eq. (3). In Fig. 4(b) the same data are shown, but instead of the phonon temperature an effective electron temperature was used to obtain a collapse down to the cryostat base temperature of 40 mK. The effective temperatures used for the various samples to obtain the collapse are shown in Fig. 4(c) for all samples. The saturation of the effective electronic temperatures at low bath temperature parallels the saturation of conductivity seen in the temperature dependence of the conductivity in Fig. 2(c).
Devices in this study are mostly arsenic $\delta$-layers, except for one low density and one high density phosphorous $\delta$-layer (see Table. 2). The strength of the WL and AA corrections are summarized in Fig. 3(b) where results from the phosphorous $\delta$-layer are indicated as squares. While arsenic dopants are expected to have stronger spin-orbit coupling than phosphorous, due to their larger mass, we find that the WL and AA effects are independent of dopant species. This indicates that spin-orbit coupling has a negligible effect on magneto-conductance.
Conclusion
In conclusion, magneto-transport measurements on $\delta$-layers of arsenic and phosphorous dopants in silicon show that the usual description of disordered metal conductivity by weak localization effects alone is insufficient in dilute and thin layers, due to significant electron-electron interaction effects. We find that when the conduction layers are sufficiently thin and dilute it becomes necessary to include the Zeeman splitting which diminishes the triplet enhancement of the conductivity PhysRevB.26.4009 ; zala2001interaction . We confirm that the effect of Zeeman splitting on the Altshuler-Aronov corrections is isotropic and combines additively with the weak-localization effect.
Acknowledgments
This project received funding from the European Research Council under the European Union’s Horizon 2020 research and innovation program, within the Hidden, Entangled and Resonating Order (HERO) project with Grant Agreement 810451. We acknowledge financial support of the Engineering and Physical Sciences Research Council (EPSRC) [grant numbers EP/R034540/1, EP/W000520/1]; and Innovate UK [75574]. J.B and P.C.C. were supported by the EPSRC Centre for Doctoral Training in Advanced Characterization of Materials (grant number EP/L015277/1), and by the Paul Scherrer Institut. P.C.C. was partially supported by Microsoft. M.M. was supported by SNSF Grant 200558.